# Mechanochemical feedback drives complex inertial dynamics in active solids

### v1

@@ Abstract
Introduction. Soft materials driven by chemical stimuli have been widely used as actuators to achieve complex tasks—locomotion, grasping, and shape-morphing—through flexible deformations [1, 2]. While promising for technologies [3], their contraction speed and power output are often limited by slow, diffusive mechanisms (thermal, fluidic, ionic, etc.) [4, 5] and damping. To overcome this limitation, recent approaches have used miniaturization [6, 7] or inertial dynamics triggered by plant-inspired elastic instabilities [8, 9, 10] and fast chemical processes (e.g., combustion) [11, 12, 13, 14] to achieve fast actuation. However, such systems passively relax without sustained input, prompting us to ask: How can soft, chemically driven materials sustain fast and continuous inertial motion without cyclic stimulation, and how complex can this motion be?
@@
where $r_{\mathrm{e}q}=(k_{0}/k_{B}T)(\mu_{A}-\mu_{B})=-2k_{0}[f^{\prime}(\chi)+C\epsilon]$ is the equilibrium reaction rate including the passive coupling to strain and $\zeta=k_{0}(h_{A}-h_{B})/k_{B}T$ is the non-equilibrium coupling to stress. Eqs. (Mechanochemical feedback drives complex inertial dynamics in active solids -3) along with initial and boundary conditions completely specify the inertial active solid model.
Figure 1: Mechanochemical coupling. (a,b) Feedback loops among reaction ($\chi$), strain ($\partial_{x}u$), strain rate ($\partial_{x}v$) in equilibrium (a) and active (b) solids. (c) Contour plot of reaction rate $\partial_{t}\chi$ (in A.U.) using the 0D model in Eq. (Mechanochemical feedback drives complex inertial dynamics in active solids ), with the dashed line marking $\partial_{t}\chi=0$ so $\chi$ grows (decays) above (below) this curve. An example trajectory is shown where a perturbation from the origin leads to a limit cycle, with four representative states illustrated in the insets.
Figure 2: Dynamics of the 1D mechanochemical system in Eq. (Mechanochemical feedback drives complex inertial dynamics in active solids ). We set $L=2\pi$, $\eta=0.16$, and $\tau_{c}=100$ along with (a) $\alpha_{u}=50,\alpha_{v}=0$, (b) $\alpha_{u}=-10,\alpha_{v}=0$, (c) $\alpha_{u}=-50,\alpha_{v}=0$, (d) $\alpha_{u}=50,\alpha_{v}=77$, (e) $\alpha_{u}=50,\alpha_{v}=80$, (f) $\alpha_{u}=50,\alpha_{v}=82$. We include a small diffusion constant $D=0.05$ to regularize chemical gradients. The initial conditions are $u(x,0)=0.01\sin(x)$, $v(x,0)=0=\chi(x,0)$. The first row shows the displacement field $u(x,t)$. The second row shows the phase space $(\partial_{x}u,v)$ trajectory at the point $x=\pi/2$.
@@
where $c=n_{0}k_{B}TC$ is the passive mechanochemical coupling, $\tau_{c}=1/[2k_{0}a_{2}+\zeta c]$ is the normalized chemical reaction timescale, $g=2k_{0}a_{4}\tau_{c}$ controls the chemical nonlinearity, and $\alpha_{u}=\tau_{c}(2k_{0}C-\zeta E)$ and $\alpha_{v}=-\tau_{c}\zeta\eta$ are the nonequilibrium feedback couplings to strain and strain-rate, respectively (Fig. 1a,b).
@@
Scaling analysis reveals the key features underlying the transitions between the qualitative behaviors exhibited in Fig. 2. In the absence of a chemical reaction ($\chi=0$), long-wavelength deformations on the scale of the system size $L$ transport momentum by elastic waves on an inertial timescale $\tau_{\mathrm{el}}\sim L\sqrt{\rho/E}$, and dissipate it through viscous stresses on a diffusive (Stokes) timescale $\tau_{\eta}\sim\rho L^{2}/\eta$. Importantly, we note that $\tau_{\eta}$ depends on inertia and is distinct from the local viscoelastic relaxation time $\tau_{ve}\sim\eta/E$. Reaction kinetics on the other hand proceeds on a chemical timescale $\sim\tau_{c}$, but nonequilibrium mechanochemical feedback can pump energy into the system at different rates. To see this, we note that a chemically induced stress $\sim c\chi$ generates a characteristic elastic strain $|\epsilon|\sim c\chi/E$ and viscous strain rate $|\dot{\epsilon}|\sim c\chi/\eta$ that drive the reaction on two new time-scales, $\tau_{1}\sim\tau_{c}E/|\alpha_{u}c|$ and $\tau_{2}\sim\tau_{c}\eta/|\alpha_{v}c|$, set by the strain and strain-rate feedback terms respectively (Fig. 1b). When activity overwhelms the rate of mechanical dissipation, i.e., $\tau_{1},\tau_{2}\lesssim\tau_{\eta},\tau_{ve}$, we expect complex dynamics to emerge (Fig. 2). Note that for the examples in Fig. 2 we choose $\eta=0.16$, so $\tau_{\eta}>\tau_{\mathrm{el}}$ and the system is mechanically underdamped.
@@
Figure 3: Dynamical phases due to nonreciprocal strain-chemistry feedback ($A_{v}=0,A_{u}<0$) in the 0D model (Eq. (Mechanochemical feedback drives complex inertial dynamics in active solids )). (a) Phase diagram ($\mathsf{Ch}=100$). (b-d) Typical trajectories in the $(U,V)$ plane for (b) the LC with $A_{u}=-11.0$, (c) one FP at $U_{0}=X_{0}=0,V_{0}=0$ with $A_{u}=-6.0$, and (d) FPs with the trajectory converging to one of them $U_{+}=-X_{+}=\sqrt{1+A_{u}}$ with $A_{u}=-0.5$. The FP is shown as red dots in (b,c), and red (unstable) and green (stable) dots in (d). We use $\mathsf{Vi}=0.10$ in all cases. (e) Dependence of the time period $T$ on $A_{u}$ for different $\mathsf{Ch}$ in the large LC region (in $\tau_{\mathrm{el}}=1$ units).
@@
Figure 4: Dynamical phases due to nonreciprocal strain-strain rate-chemistry feedback ($A_{v}>0,A_{u}>0$) in the 0D model (Eq. (Mechanochemical feedback drives complex inertial dynamics in active solids )). (a) Phase diagram ($\mathrm{Ch}=100$, $A_{u}=50$). (b-f) Typical trajectories in the $(U,V)$ plane for the FPs with the trajectory converging to (b) one FP $U_{+}=-X_{+}=\sqrt{1+A_{u}}$ for $A_{v}=60.5$, (c) an LC around one unstable FP for $A_{v}=68.0$, (d) period doubling for $A_{v}=78.5$, (e) chaos for $A_{v}=84.0$, and (f) the large LC conclusing all FPs for $A_{v}=106.0$. In all cases, we set $\mathsf{Vi}=0.16$ and $A_{u}=50.0$. (g) Bifurcation diagram showing the route from LC around the FP $U_{\pm}=-X_{\pm}=\sqrt{1+A_{u}}$ to chaos as a function of the parameter $A_{v}$. Here, $u_{\mathrm{m}}$ represents the local maximum value of $U$. (h) Dependence of the time period $T$ on $A_{v}$ at $\mathsf{Vi}=0.16$ in the large LC region, approaching $2\pi$ as $A_{v}$ increases (in $\tau_{\mathrm{el}}=1$ units).
@@
At $A_{v}=0$, nonreciprocity between strain and reaction arises for $A_{u}<0$. Upon decreasing $A_{u}$, a linear stability analysis reveals first a pitchfork bifurcation where 3 FPs (bistable) transition to 1 FP (monostable) and then an oscillatory Hopf bifurcation occurs at a second threshold when $A_{u}\approx-\mathsf{Vi}\;\mathsf{Ch}$ [71] (Fig. 3). The analytical bifurcation thresholds (derived in the SM [57]) are confirmed via numerical continuation using MATCONT and shown in Fig. 3a. This Hopf bifurcation corresponds to $\tau_{1}\lesssim\tau_{\eta}$, i.e., when the strain feedback actively pumps energy faster than viscous stresses can dissipate it, and it crucially depends on inertia as $\tau_{\eta}\sim\rho L^{2}/\eta$.
@@
Near the Hopf bifurcation, the imaginary part of the unstable eigenvalue yields the time period of the LC to be $T=2\pi\tau_{\mathrm{el}}/\sqrt{1-(\mathsf{Vi}/\mathsf{Ch})}\simeq 2\pi\tau_{\mathrm{el}}$ for $\mathsf{Ch}\gg 1$ (dashed line in Fig. 3e): fast oscillations at the inertial time scale, distinct from previous examples of spontaneous oscillations in overdamped active solids [72, 19, 27, 16, 17]. Interestingly, this instability and LC behavior persists even in the mechanically overdamped regime ($\mathsf{Vi}>1$, Fig. 3) with a period controlled by the same $\tau_{el}$, as long as nonreciprocal feedback overwhelms the mechanical dissipation rate (see more details in the SM [57]).
@@
Now we consider $A_{u}>0,A_{v}>0$, where strain rate steers the activity (Fig. 4). By choosing $A_{u}>0$, we always have three FPs, with the FP at the origin ($U_{0},X_{0}$) always unstable. A linear stability analysis shows that upon increasing $A_{v}$, both nontrivial FPs $(U_{\pm},X_{\pm})$ lose stability via a Hopf bifurcation when $A_{v}\gtrsim\mathsf{Ch}$ (valid in the limit of slow kinetics, i.e., $\mathsf{Ch}\gg\mathsf{Vi},A_{u}$, see SM [57] for the full expression). This analytical bifurcation threshold (white dashed line in Fig. 4(a)) is confirmed by numerical analysis. In terms of physical time scales, this bifurcation condition corresponds to $\tau_{2}<\tau_{ve}$, i.e., the strain rate feedback time is smaller than the local viscoelastic relaxation time. Near onset, the small LC has a slow oscillation time period $T_{\mathrm{small}}\approx\pi\tau_{\mathrm{el}}\sqrt{2\mathsf{Vi}\,\mathsf{Ch}/(1+A_{u})}\sim\sqrt{\eta\tau_{c}/E}$ governed by dissipative processes. Note that these timescales do not rely on either spatial gradients or inertia, and the resulting instability is smoothly connected to previously studied instabilities in overdamped active solids [72, 19, 27, 16, 17].
@@
At larger $A_{v}$, more complex dynamics arises, enabled by the three-dimensional phase space ($U,V,X$) including inertia. We find a period doubling cascade for the pair of LCs, which collide into a chaotic strange attractor (Fig. 4e), reminiscent of the classical Lorenz and Rossler system [73]. Chaos develops in a large parameter regime in the $A_{v}-\mathsf{Vi}$ space (Fig. 4a), and is interspersed by small islands of periodic orbits (Fig. 4g and SM [57] for Lyapunov exponent analysis). The phase boundaries separating chaos from limit cycles in Fig. 4(a) are obtained by analyzing phase-space trajectories obtained by direct numerical simulations. For even larger values of $A_{v}$, eventually, the system settles into a globally attracting large LC that encloses all three FPs (Fig. 4f). While the small LCs have slow oscillations, we numerically find that the large LC exhibits fast inertial oscillations with a time period approaching $T\approx 2\pi\tau_{\mathrm{el}}$ as $A_{v}\to\infty$ (Fig. 4h). This explains and recapitulates the complex behavior observed in the 1D model (Fig. 2).
@@
Conclusion. By using nonequilibrium feedback to couple chemical reactions with the mechanics of a soft solid, we have shown how active driving can sustain autonomous deformations, fast oscillations, and nonlinear dynamics. These results highlight the importance of inertia in active mechanochemical solids and provide a design principle for generating high quality ultrafast oscillators even in dissipative, soft solids. That feedback can also trigger chaos in active solids suggests a simple way to generate complex “twitching” movements, reminiscent of rudimentary behaviors in soft-bodied organisms [74, 15]. Our work paves the way to explore a new regime of chemically active solids in previously inaccessible high frequency scales, and provides a scalable and efficient platform to design autonomous soft materials from homogeneous media, without the need for complex microstructures or discretely embedded active particles.

### v2

@@ Abstract
Introduction. Soft materials driven by chemical stimuli have been widely used as actuators to achieve complex tasks—locomotion, grasping, and shape-morphing—through flexible deformations [1, 2]. While promising for technologies [3], their contraction speed and power output are often limited by slow, diffusive mechanisms (thermal, fluidic, ionic, etc.) and damping [4, 5]. To overcome this limitation, recent approaches have used miniaturization [6, 7] or inertial dynamics triggered by plant-inspired elastic instabilities [8, 9, 10] and fast chemical processes (e.g., combustion) [11, 12, 13, 14] to achieve fast actuation. However, such systems passively relax without sustained input, prompting us to ask: How can soft, chemically driven materials sustain fast and continuous inertial motion without cyclic stimulation, and how complex can this motion be?
@@
where $r_{\mathrm{e}q}=(k_{0}/k_{B}T){(\mu^{p}_{A}-\mu^{p}_{B})}=-2k_{0}[f^{\prime}(\chi)+C\epsilon]$ is the equilibrium reaction rate including the passive coupling to strain and $\zeta=k_{0}(h_{A}-h_{B})/k_{B}T$ is the non-equilibrium coupling to stress. Note that for the system to be passive, we need $\zeta=0$, which invalidates coupling to strain rate. Eqs. (Mechanochemical feedback drives complex inertial dynamics in active solids-3) along with initial and boundary conditions completely specify the inertial active solid model.
Figure 1: Mechanochemical coupling. (a,b) Feedback loops among reaction ($\chi$), strain ($\partial_{x}u$), strain rate ($\partial_{x}v$) in passive (a) and active (b) solids. In the passive limit (a), detailed balance enforces vanishing strain-rate coupling, and a reciprocal strain-reaction coupling ($k_{B}TC$). When active (b), the feedback drives nonequilibrium fluxes on independent time scales $\tau_{1,2}$. (c) Contour plot of reaction rate $\partial_{t}\chi$ (in A.U.) using the 0D model in Eq. (Mechanochemical feedback drives complex inertial dynamics in active solids), with the dashed line marking $\partial_{t}\chi=0$ so $\chi$ grows (decays) above (below) this curve. An example trajectory is shown where a perturbation from the origin leads to a limit cycle, with four representative states illustrated as insets, where the background shading refers to concentration $\chi$. Starting with the bottom left corner, increased strain breaks stress-sensitive crosslinkers, which drives $\chi$ to increase, followed by decrease in strain, and recovery back to the first state.
Figure 2: Dynamics of the 1D mechanochemical system in Eq. (Mechanochemical feedback drives complex inertial dynamics in active solids). We set $L=2\pi$, $\eta=0.16$, and $\tau_{c}=100$ along with (a) $\alpha_{u}=50,\alpha_{v}=0$, (b) $\alpha_{u}=-10,\alpha_{v}=0$, (c) $\alpha_{u}=-50,\alpha_{v}=0$, (d) $\alpha_{u}=50,\alpha_{v}=77$, (e) $\alpha_{u}=50,\alpha_{v}=80$, (f) $\alpha_{u}=50,\alpha_{v}=82$. We include a small diffusion constant $D=0.05$ to regularize chemical gradients. The initial conditions are $u(x,0)=0.01\sin(x)$, $v(x,0)=0=\chi(x,0)$ The first row shows the displacement field $u(x,t)$. The second row shows the phase space $({\epsilon=\partial_{x}u},v)$ trajectory at the point $x=9\pi/25$.
@@
where $c=n_{0}k_{B}TC$ is the passive chemically induced stress scale, $\tau_{c}=1/[2k_{0}a_{2}+\zeta c]$ is the effective chemical reaction timescale, combining the equilibrium rate ($2k_{0}a_{2}$) and the active feedback rate ($\zeta c$[57], $g=2k_{0}a_{4}\tau_{c}$ controls the chemical nonlinearity, and $\alpha_{u}=\tau_{c}(2k_{0}C-\zeta E)$ and $\alpha_{v}=-\tau_{c}\zeta\eta$ are the nonequilibrium feedback couplings to strain and strain-rate, respectively (Fig. 1a,b).
@@
Scaling analysis reveals the key features underlying the transitions between the qualitative behaviors exhibited in Fig. 2. In the absence of a chemical reaction ($\chi=0$), long-wavelength deformations on the scale of the system size $L$ transport momentum by elastic waves on an inertial timescale $\tau_{\mathrm{el}}\sim L\sqrt{\rho/E}$, and dissipate it through viscous stresses on a diffusive (Stokes) timescale $\tau_{\eta}\sim\rho L^{2}/\eta$. Importantly, we note that $\tau_{\eta}$ depends on inertia and is distinct from the local viscoelastic relaxation time $\tau_{ve}\sim\eta/E$. The bare reaction kinetics on the other hand proceeds on a chemical timescale $\sim\tau_{c}$, but nonequilibrium mechanochemical feedback can pump energy into the system at independent rates. To see this, we consider the time taken for a minimal feedback cycle (Fig. 1b,c). A chemically induced stress $\sim c\chi$ generates a characteristic elastic strain $|\epsilon|\sim c\chi/E$ or viscous strain rate $|\dot{\epsilon}|\sim c\chi/\eta$ that can in turn drive the reaction via $\tau_{c}\dot{\chi}\sim\alpha_{u}\epsilon$ or $\tau_{c}\dot{\chi}\sim\alpha_{v}|\dot{\epsilon}|$, completing the cycle. This gives two new time-scales, $\tau_{1}\sim\tau_{c}E/|\alpha_{u}c|$ and $\tau_{2}\sim\tau_{c}\eta/|\alpha_{v}c|$, set by the strain and strain-rate feedback terms respectively (Fig. 1b). When activity overwhelms the rate of mechanical dissipation, i.e., $\tau_{1},\tau_{2}\lesssim\tau_{\eta},\tau_{ve}$, we expect complex dynamics to emerge (Fig. 2). Note that for the examples in Fig. 2 we choose $\eta=0.16$, so $\tau_{\eta}>\tau_{\mathrm{el}}$ and the system is mechanically underdamped.
@@
Figure 3: Dynamical phases due to nonreciprocal strain-chemistry feedback ($A_{v}=0,A_{u}<0$) in the 0D model (Eq. (Mechanochemical feedback drives complex inertial dynamics in active solids)). (a) Phase diagram ($\mathsf{Ch}=100$). (b-d) Typical trajectories in the $(U,V)$ plane for (b) the outer (larger) LC with $\left(\mathsf{Vi},A_{u}\right)=\left(0.70,-130.0\right)$ (underdamped regime) and the inner (smaller) LC with $\left(\mathsf{Vi},A_{u}\right)=\left(1.2,-140.0\right)$ (overdamped regime), (c) one FP at $U_{0}=X_{0}=0,V_{0}=0$ with $\left(\mathsf{Vi},A_{u}\right)=\left(0.70,-20.0\right)$, and (d) FPs with the trajectory converging to one of them $U_{+}=-X_{+}=\sqrt{1+A_{u}}$ with $\left(\mathsf{Vi},A_{u}\right)=\left(0.70,-0.5\right)$. The FPs are shown as red (unstable) and green (stable) dots in (b-d), with the arbitrarily chosen initial conditions denoted by the black star. In panels (b-d), different symbols adjacent to the panel index represent the corresponding points in the phase diagram in (a). (e) Dependence of the time period $T$ on $A_{u}$ for different $\mathsf{Ch}$ in the LC region (in $\tau_{\mathrm{el}}=1$ units).
@@
Figure 4: Dynamical phases due to nonreciprocal strain-strain rate-chemistry feedback ($A_{v}>0,A_{u}>0$) in the 0D model (Eq. (Mechanochemical feedback drives complex inertial dynamics in active solids)). (a) Phase diagram ($\mathrm{Ch}=100$, $A_{u}=50$). (b-f) Typical trajectories in the $(U,V)$ plane for the FPs with the trajectory converging to (b) one FP $U_{+}=-X_{+}=\sqrt{1+A_{u}}$ for $A_{v}=60.5$, (c) an LC around one unstable FP for $A_{v}=68.0$, (d) period doubling for $A_{v}=78.5$, (e) chaos for $A_{v}=84.0$, and (f) the large LC enclosing all FPs for $A_{v}=106.0$. In all cases, we set $\mathsf{Vi}=0.16$ and $A_{u}=50.0$. The black stars in (b-f) denote the arbitrarily chosen initial condition. (g) Bifurcation diagram showing the route from LC around the FP $U_{\pm}=-X_{\pm}=\sqrt{1+A_{u}}$ to chaos as a function of the parameter $A_{v}$. Here, $u_{\mathrm{m}}$ represents the local maximum value of $U$ [57]. (h) Dependence of the time period $T$ on $A_{v}$ at $\mathsf{Vi}=0.16$ in the large LC region, approaching $2\pi$ as $A_{v}$ increases (in $\tau_{\mathrm{el}}=1$ units).
@@
At $A_{v}=0$, nonreciprocity between strain and reaction arises for $A_{u}<0$. Upon decreasing $A_{u}$, a linear stability analysis reveals first a pitchfork bifurcation where 3 FPs (bistable) transition to 1 FP (monostable) and then an oscillatory Hopf bifurcation occurs at a second threshold when $A_{u}\approx-\mathsf{Vi}\;\mathsf{Ch}$ [71] (Fig. 3). Figure 3 shows typical trajectories in the $(U,V)$ plane for limit cycles (b) and FPs (c-d). The analytical bifurcation thresholds (derived in the SM [57]) are confirmed via numerical continuation using MATCONT and shown in Fig. 3a. This Hopf bifurcation corresponds to $\tau_{1}\lesssim\tau_{\eta}$, i.e., when the strain feedback actively pumps energy faster than viscous stresses can dissipate it, and it crucially depends on inertia as $\tau_{\eta}\sim\rho L^{2}/\eta$.
@@
Near the Hopf bifurcation, the imaginary part of the unstable eigenvalue yields the time period of the LC to be $T=2\pi\tau_{\mathrm{el}}/\sqrt{1-(\mathsf{Vi}/\mathsf{Ch})}\simeq 2\pi\tau_{\mathrm{el}}$ for $\mathsf{Ch}\gg 1$ (dashed line in Fig. 3e): fast oscillations at the inertial time scale, distinct from previous examples of spontaneous oscillations in overdamped active solids [72, 19, 27, 16, 17]. Interestingly, this instability and LC behavior persists even in the mechanically overdamped regime ($\mathsf{Vi}>1$, see SM [57] for more details) with a period controlled by the same $\tau_{\mathrm{el}}$, as long as nonreciprocal feedback overwhelms the mechanical dissipation rate (Fig. 3b,e).
@@
Now we consider $A_{u}>0,A_{v}>0$, where strain rate steers the activity (Fig. 4). By choosing $A_{u}>0$, we have three FPs, with the FP at the origin ($U_{0},X_{0}$) always unstable. A linear stability analysis shows that upon increasing $A_{v}$, both nontrivial FPs $(U_{\pm},X_{\pm})$ lose stability via a Hopf bifurcation when $A_{v}\gtrsim\mathsf{Ch}$ (valid in the limit of slow kinetics, i.e., $\mathsf{Ch}\gg\mathsf{Vi},A_{u}$, see SM [57] for the full expression). This analytical bifurcation threshold (white dashed line in Fig. 4(a)) is confirmed by numerical analysis. In terms of physical time scales, this bifurcation condition corresponds to $\tau_{2}<\tau_{ve}$, i.e., the strain rate feedback time is smaller than the local viscoelastic relaxation time. Near onset, the small LC has a slow oscillation time period $T_{\mathrm{small}}\approx\pi\tau_{\mathrm{el}}\sqrt{2\mathsf{Vi}\,\mathsf{Ch}/(1+A_{u})}\sim\sqrt{\eta\tau_{c}/E}$ governed by dissipative processes. Note that these timescales do not rely on inertia, and the resulting instability is smoothly connected to previously studied instabilities in overdamped active solids [72, 19, 27, 16, 17].
@@
Enabled by the three-dimensionality of phase space ($U,V,X$) including inertia, at large $A_{v}$, chaos arises. We find a period doubling (Fig. 4d) cascade for the pair of LCs, which collide into a chaotic strange attractor (Fig. 4e), reminiscent of the classical Lorenz and Rossler system [73]. Chaos develops in a large parameter regime in the $A_{v}-\mathsf{Vi}$ space (Fig. 4a), and is interspersed by small islands of periodic orbits (Fig. 4g and SM [57] for Lyapunov exponent analysis). The phase boundaries separating chaos from limit cycles in Fig. 4(a) are obtained by analyzing phase-space trajectories obtained by direct numerical simulations. For even larger values of $A_{v}$, eventually, the system settles into a globally attracting large LC that encloses all three FPs (Fig. 4f). While the small LCs have slow oscillations, we numerically find that the large LC exhibits fast inertial oscillations with a time period approaching $T\approx 2\pi\tau_{\mathrm{el}}$ as $A_{v}\to\infty$ (Fig. 4h). This explains and recapitulates the complex behavior observed in the 1D model (Fig. 2). Whether such chaotic and fast dynamics observed for $A_{v}>0$ survives when $\mathsf{Vi}\gtrsim 1$ is left for future work.
@@
Conclusion. By using nonequilibrium feedback to couple chemical reactions with the mechanics of a soft solid, we have shown how active driving can sustain autonomous deformations, fast oscillations, and complex dynamics. These results highlight the importance of inertia in active mechanochemical solids and provide a design principle for generating high quality ultrafast oscillators even in dissipative, soft solids. That feedback can also trigger chaos in active solids suggests a simple way to generate complex “twitching” movements, reminiscent of rudimentary behaviors in soft-bodied organisms [74, 15]. Our work paves the way to explore a new regime of chemically active solids in previously inaccessible high frequency scales, and provides a scalable and efficient platform to design autonomous soft materials from homogeneous media, without the need for complex microstructures or discretely embedded active particles.
[…]